OpenStudy (anonymous):

write the equation of the line which passes through (2,-3) and is perpendicular to y=4x+7 in standard form

OpenStudy (chaise):

Perpdinduclar lines' gradient multiply to give -1 y = 4x+7 The gradient of the perpindcular line will be "-1/4" m = -1/4 y=mx+b you know the line runs through points (2,-3). You can substitute these values into your formula along with m to solve for b (y intercept) -3 = -1/4x3 +b -3 = -0.75+b b = -2.25 y=mx+b y = -1/4-2.25 Hope this helped. Let me know if you didn't understand how I got this answer :)

OpenStudy (chaise):

Where I multiplied with a 3 it should be a 2, my bad. This will give a answer of -2.5 ): Let me rewrite this. -3 = -1/4x2 +b -3 = -0.5+b b = -2.5

OpenStudy (anonymous):

is this like y =mx+b